. Henry invested $12,000 in a new restaurant. When the restaurant was sold two years
later, he received $27,000. Find his average annual return by solving the equation 12,000(1 + r)^2 =
27,000.

Answer by checkley77(12844)

12,000(1+r)^2=27,000.
(1+r)^2=27,000/12,000
(1+r)^2=2.25 take the sqrt of both sides.
1+r=1.5
r=1.5-1
r=.5 or 50% annual return.